A Saleslady sold 30 bags at an average price of Php350. She got Php6,800 for 20 of them. At what price did she sell each of the last 10 bags?
Answer by Cromlix(4381)

Average price of Php350 for 30 bags
Total for 30 bags = 30 * 350 = Php10500
She got Php6,800 for 20 bags
The 10 bags left must have cost
Php10500 - Php6,800 = Php3700
Php3700 / 10 = Php370
The lady sold the last 10
bags for Php370 each.
Hope this helps.
:-)